An electron with KE of 2keV collided with a positron at rest creating two photons that have equal energy. Show that the energy of each photon is 8.2{times 10}-14J
1
Expert's answer
2018-04-27T14:36:07-0400
Form the conservation of energy:
mc^2+mc^2+KE=2E
E=1/2 KE+mc^2=1/2 2+511=512 keV.
E=512(1000)1.6∙〖10〗^(-19) J=8.2∙〖10〗^(-14) J